write an equation:
slope = 1/3 passing through (-1,5)

Answer by Alan3354(69443)

write an equation:
slope = 1/3 passing through (-1,5)
---------------
Use y = mx+b and solve for b
5 = (1/3)*(-1) + b
b = 16/3
--> y = (1/3)x + 16/3
or x - 3y = -16